OpenStudy (nightwing255):

Check all the ordered pairs that satisfy the equation below y=3/4x check all that apply? A. (16,12 ) B. (5, 15) C. (21,14) D. (28, 21) E. (12,4/3 ) F. (18, 24)

jimthompson5910 (jim_thompson5910):

How far did you get with this problem?

OpenStudy (nightwing255):

Not so far

jimthompson5910 (jim_thompson5910):

what you do for each answer choice is plug in the given x and y values

jimthompson5910 (jim_thompson5910):

so for instance, in choice A we have x = 16 and y = 12 so, y = (3/4)*x 12 = (3/4)*16 12 = 12 ... true equation since the last equation is true, this means that (16,12) is a point on the line. Hopefully this makes sense?

OpenStudy (nightwing255):

It makes sense, thank you

jimthompson5910 (jim_thompson5910):

you're welcome. Let me know if you get stuck anywhere

OpenStudy (nightwing255):

I'm actually good right now thanks

jimthompson5910 (jim_thompson5910):

sure thing

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)